OpenStudy (anonymous):

If a line crosses the y-axis at (0, -3) and has a slope of 3, what is the equation of the line?

OpenStudy (anonymous):

slope intercept form is y = mx + b

OpenStudy (anonymous):

m being your slope and b being your y-intercept

OpenStudy (anonymous):

so your slope is 3 as it says and it also says that line crosses through (0, -3) which is your y-intercept so the equation is y = 3x -3 :)

OpenStudy (anonymous):

what dont you understand michael?

OpenStudy (anonymous):

-3=3+0

OpenStudy (anonymous):

@micheal05 what dont you understand?

OpenStudy (anonymous):

how to do it

OpenStudy (anonymous):

how to make the equation?

OpenStudy (anonymous):

yahhh

OpenStudy (anonymous):

y = mx + b

OpenStudy (anonymous):

can you tell me what the m and b stand for

OpenStudy (anonymous):

ok ik that but like idk how to put the numbers into that

OpenStudy (anonymous):

ok so the m in the equation is your slope so if you have like a slope of 6 then your equation would look like y = 6x

OpenStudy (anonymous):

and your b is your y intercept so if you had a y-intercept of 5 then your entire equation would look like y = 6x + 5

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

np